OpenStudy (anonymous):

Find all of the points on the line y = 2x + 5 which are 4 units from the point (−1, 7).

OpenStudy (anonymous):

if the point is \((a,b)\) the you know it looks like \((a, 2a+5)\) and so the square of the distance between that point and \((-1,7)\) will be \[(a+1)^2+(2a+5-7)^2=16\] or \[(a+1)^2+(2a-2)^2=16\] solve the quadratic equation for \(a\)

OpenStudy (whpalmer4):

You know the formula for the distance between two points. \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] Use your known point as one of the points. \[\sqrt{(x+1)^2 + (2x+5-7)^2} = 4\]\[(x+1)^2 + (2x-2)^2 = 16\]\[x^2+2x+2 + 4x^2 - 8x + 4 = 16\]\[5x^2 -6x -10 = 0\]Solving that for \(x\) will give you the x values of the points, and the y-values you'll find via \(y = 2x+5\)

OpenStudy (anonymous):

\[(a+1)^2+(2a-2)^2=16\] \[a^2+2a+1+4a^2-8a+4=16\] \[5a^2-6a-11=0\] solve for \(a\)

OpenStudy (whpalmer4):

Oops, yes, how'd I do that? Thanks @satellite73 !

OpenStudy (anonymous):

Thank you guys :)